OpenStudy (anonymous):

Please, check if I'm correct. 5).Complete the square: 0 = - 5x^2 + 10x - 10 (A) (x - 1)^2 = -5 (B) (x - 1)^2 = 5 (C) (x - 1)^2 = - 1******* 7). If the discriminant is positive-- if b^2 -4ac > 0 --then the quadratic equation (A) has two solutions******** (B) has one solution (C) has no solutions

OpenStudy (king):

7.A

OpenStudy (anonymous):

You're right! 5C and 7A :) - 5x^2 + 10x = 10 -5 ( x^2 - 2x) =10 -5 [ (x -1)^2 - 1 ] = 10 (x -1)^2 - 1 = -2 (x -1)^2 = -1

OpenStudy (anonymous):

That's the right answer. Thanks.